# Premise Order Matters in Reasoning with Large Language Models

### v1

@@ 1 Introduction
We conduct a systematic study on the premise order effect using a variety of SoTA LLMs, including GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), PaLM 2-L (Google, 2023), and Gemini Pro (Gemini, 2023). Our primary focus is deductive reasoning, and we benchmark all LLMs on problems that only involve modus ponens (if P then Q; P; therefore Q), where all LLMs in our evaluation at least achieve decent performance with a small number of premises. We show that the accuracy decrease caused by different ordering can be more than 30%. The ordering effect is further amplified when irrelevant premises (i.e., premises that are not needed to derive a conclusion) are presented in the prompt. Figure 1 illustrates a failure case, where all LLMs fail to generate the proof after changing the order of relevant rules. Interestingly, while all LLMs perform best when the premise order follows the ground truth proof, they reveal different preferences on other alternative orderings. Specifically, compared to randomly ordering the premises, GPT-4-turbo and GPT-3.5-turbo generally achieve better performance when the premise order is exactly the reverse of the ground truth proof, which enables LLMs to perform derivation via backward chaining. On the other hand, PaLM 2-L generally achieves the worst performance with such a reversed order.
@@ 2.1 Logical Reasoning
Prior work has revealed the weaknesses of LLMs in logical reasoning (Han et al., 2022; Xu et al., 2023; Saparov et al., 2023; Saparov and He, 2022; Wan et al., 2024), especially when the proof is long and requires the knowledge of multiple deduction theorems. To isolate the effect of premise orders, we focus on a confined problem space adapted from SimpleLogic (Zhang et al., 2022), which only includes propositional logic problems with definite clauses. Specifically, each problem includes: (1) a set of facts $A_{1}$$\ldots$, $A_{n}$ that hold true; (2) a set of rules of the form “If $X$, then $Y$”, “If $X_{0}$ and $X_{1}$, then $Y$”, or “If $X_{0}$ and $X_{1}$ and $X_{2}$, then $Y$”; and (3) a conclusion “$C$ is True” to be proved. As opposed to SimpleLogic — which formulates the problem as a binary classification task (i.e., indicate whether the conclusion is True or False) — in our benchmark, every problem has a ground-truth label of True, and we consider the prediction to be correct only when the generated proof is completely valid. With these strict criteria, the LLM is required to produce the step-by-step deduction that leads to the conclusion, and any hallucination of non-existent facts and rules is considered erroneous.
@@ 3.1 Experimental Setup
We evaluate the premise ordering effect on GPT-4-turbo, GPT-3.5-turbo, PaLM 2-L and Gemini Pro. We perform the greedy decoding with the temperature 0, and apply the zero-shot prompting in all experiments. On R-GSM, the model input only contains the problem description without additional instructions. For logical reasoning, as shown in Figure 1, we add an instruction in the prompt to ask for a derivation that specifies which premise is used in each step.
@@ 3.2 Logical Reasoning
Figure 3: Logical reasoning without distracting rules. See Table 5 in Appendix D for accuracy numbers.
Figure 4: Logical reasoning with distracting rules. See Tables 6 and 7 for accuracy numbers.
@@
Figure 6: Results on different $\tau$ with distracting rules. See Tables 9 and 10 for accuracy numbers.
Figure 3 presents the results with different numbers of relevant rules included in ground truth proofs, where the problem does not contain distracting rules, and the shuffled accuracy is the aggregation of results with $\tau$ = 0.5, 0 and -0.5. Across different LLMs, the forward order consistently achieves the best performance, which aligns with the human preference. The performance drop caused by alternative orderings becomes more significant when the number of rules increases. Meanwhile, models with weaker reasoning capabilities are also more sensitive to different premise orders. Specifically, while the accuracy decrease of GPT-4-turbo and PaLM 2-L is up to $20-30\%$, with Gemini-Pro and GPT-3.5-turbo, changing the premise order from the forward order can degrade the accuracy from over $65\%$ to below $25\%$, with an accuracy decrease of more than $40\%$.
Breakdown on different premise orders. We present the results of fine-grained breakdown on premise ordering in Figure 5, where the orders are categorized based on Kendall tau distance $\tau$ as described in Section 2.1. Interestingly, while the top preference of all LLMs is the forward order, their preferences on other orders are not alike. Specifically, GPT-4-turbo generally prefers the backward order over other orders, and the overall performance decreases with a smaller absolute value of $\tau$. This observation is also consistent with the human reasoning pattern, as backward chaining is another well-established inference method. On the other hand, PaLM 2-L generally performs the worst with the backward order. With the decrease of $\tau$ (i.e., the premise order deviates more from the forward order), the accuracy drops. The preferences of Gemini Pro and GPT-3.5-turbo are less consistent, still they prefer the backward order more often than other non-forward premise orders.
@@ 3.3 R-GSM for Mathematical Reasoning
Figure 7: R-GSM results with different numbers of reasoning steps in the ground truth. See Table 11 in Appendix E for accuracy numbers.
@@
Breakdown of problem complexity. Figures 7 and 8 present the breakdown results on different number of reasoning steps and different number of problem sentences, respectively. Unsurprisingly, across all LLMs, the proof accuracy suffers on problems that require more reasoning steps and contain a greater number of sentences. Overall, the gap between the accuracies on initial and rewritten problems is more significant with more reasoning steps and longer problems for both GPT-4-turbo and Gemini Pro, while the gap remains similar across different numbers of reasoning steps and problem lengths for PaLM 2-L and GPT-3.5-turbo.
@@ 4 Related Work
Failure modes of LLMs. The premise order effect in this work is connected to several failure modes of LLMs in the literature, including the reversal curse (Berglund et al., 2023), distractibility (Shi et al., 2023), and limited capability of logical reasoning (Han et al., 2022; Xu et al., 2023; Saparov et al., 2023; Saparov and He, 2022; Wan et al., 2024; Zhu et al., 2023). Specifically, Shi et al. (2023) show that including irrelevant context in the problem statement leads to a considerable performance drop on GSM8K and other reasoning benchmarks, revealing that LLMs are distractible. This finding is in-line with our evaluation on logical reasoning, where we observe that adding irrelevant rules not only degrades the overall logical reasoning performance, but also escalates the premise order effect. The Reversal Curse (Berglund et al., 2023) unveils another perspective of the order effect, where they show that an LLM that recognizes “A is B” does not necessarily learn that “B is A.” While their work studies the order effect between two entities within a single factual statement, our work focuses on reasoning problems with multiple premises, without restrictions on the number of (or relationship between) entities. In particular, for logical reasoning, we demonstrate that random permutations of premises often result in worse accuracy than the purely backward order.
@@
Order effect of language models. Some prior works show that language models are able to understand permuted texts to some extent, i.e., after a random permutation of words, models usually preserve a reasonable performance (Abdou et al., 2022; Sinha et al., 2020). Moreover,  Cao et al. (2023) shows that even when a large fraction of words are scrambled, GPT-4 still achieves decent performance on several reasoning benchmarks. In contrast to permuted texts in these works that are typically unnatural and nonsensical, our premise order permutations do not alter the semantic meaning and remain syntactically valid (we manually verify this). Nevertheless, we demonstrate that LLM reasoning performance is highly brittle to the ordering of the premises.
@@ Appendix C R-GSM Examples
Appendix D Full Results for Logical Reasoning
Tables 5 and 8 present the accuracy numbers for Figures 3 and 5, which are results on different numbers of relevant rules without distracting rules.
Tables 6 and 9 present the accuracy numbers for Figures 4 and 6 with 5 distracting rules.
Tables 7 and 10 present the accuracy numbers for Figures 4 and 6 with 10 distracting rules.
# Rules Order Acc
4 Forward 100%
Backward 100%
Shuffled 99.8%
5 Forward 100%
Backward 100%
Shuffled 99.5%
6 Forward 100%
Backward 100%
Shuffled 99.3%
7 Forward 99.5%
Backward 98.5%
Shuffled 97.8%
8 Forward 99.5%
Backward 98.5%
Shuffled 95.8%
9 Forward 99.0%
Backward 95.5%
Shuffled 95.3%
10 Forward 99.0%
Backward 97.0%
Shuffled 91.0%
11 Forward 98.5%
Backward 95.5%
Shuffled 91.5%
12 Forward 98.0%
Backward 86.5%
Shuffled 84.2%
(a) GPT-4-turbo.
@@ 6 Forward 100%
Appendix E Full Results on R-GSM
Tables 11 and 12 present the accuracy numbers for Figures 7 and 8, which are breakdown results on R-GSM problems with different numbers of reasoning steps and different numbers of sentences in the problem description respectively.

### v3

@@ 1 Introduction
We conduct a systematic study on the premise order effect using a variety of SoTA LLMs, including GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), PaLM 2-L (Google, 2023), and Gemini 1.0 Pro (Gemini, 2023). Our primary focus is deductive reasoning, and we benchmark all LLMs on problems that only involve modus ponens (if P then Q; P; therefore Q), where all LLMs in our evaluation at least achieve decent performance with a small number of premises. We show that the accuracy decrease caused by different ordering can be more than 30%. The ordering effect is further amplified when irrelevant premises (i.e., premises that are not needed to derive a conclusion) are presented in the prompt. Figure 1 illustrates a failure case, where all LLMs fail to generate the proof after changing the order of relevant rules. Interestingly, while all LLMs perform best when the premise order follows the ground truth proof, they reveal different preferences on other alternative orderings. Specifically, compared to randomly ordering the premises, GPT-4-turbo and GPT-3.5-turbo generally achieve better performance when the premise order is exactly the reverse of the ground truth proof, which enables LLMs to perform derivation via backward chaining. On the other hand, PaLM 2-L generally achieves the worst performance with such a reversed order.
@@ 2.1 Logical Reasoning
Prior work has revealed the weaknesses of LLMs in logical reasoning (Han et al., 2022; Xu et al., 2023; Saparov et al., 2023; Saparov and He, 2022; Wan et al., 2024; Yan et al., 2023), especially when the proof is long and requires the knowledge of multiple deduction theorems. To isolate the effect of premise orders, we focus on a confined problem space adapted from SimpleLogic (Zhang et al., 2022), which only includes propositional logic problems with definite clauses. Specifically, each problem includes: (1) a set of facts $A_{1}$$\ldots$, $A_{n}$ that hold true; (2) a set of rules of the form “If $X$, then $Y$”, “If $X_{0}$ and $X_{1}$, then $Y$”, or “If $X_{0}$ and $X_{1}$ and $X_{2}$, then $Y$”; and (3) a conclusion “$C$ is True” to be proved. As opposed to SimpleLogic — which formulates the problem as a binary classification task (i.e., indicate whether the conclusion is True or False) — in our benchmark, every problem has a ground-truth label of True, and we consider the prediction to be correct only when the generated proof is completely valid. With these strict criteria, the LLM is required to produce the step-by-step deduction that leads to the conclusion, and any hallucination of non-existent facts and rules is considered erroneous.
@@ 3.1 Experimental Setup
We evaluate the premise ordering effect on GPT-4-turbo, GPT-3.5-turbo, PaLM 2-L and Gemini 1.0 Pro. We perform the greedy decoding with the temperature 0, and apply the zero-shot prompting in all experiments. On R-GSM, the model input only contains the problem description without additional instructions. For logical reasoning, as shown in Figure 1, we add an instruction in the prompt to ask for a derivation that specifies which premise is used in each step.
@@ 3.2 Logical Reasoning
Figure 3: Logical reasoning without distracting rules. See Table 6 in Appendix E for accuracy numbers.
Figure 4: Logical reasoning with distracting rules. See Tables 7 and 8 for accuracy numbers.
@@
Figure 6: Results on different $\tau$ with distracting rules. See Tables 10 and 11 for accuracy numbers.
Figure 3 presents the results with different numbers of relevant rules included in ground truth proofs, where the problem does not contain distracting rules, and the shuffled accuracy is the aggregation of results with $\tau$ = 0.5, 0 and -0.5. Across different LLMs, the forward order consistently achieves the best performance, which aligns with the human preference. The performance drop caused by alternative orderings becomes more significant when the number of rules increases. Meanwhile, models with weaker reasoning capabilities are also more sensitive to different premise orders. Specifically, while the accuracy decrease of GPT-4-turbo and PaLM 2-L is up to $20-30\%$, with Gemini 1.0 Pro and GPT-3.5-turbo, changing the premise order from the forward order can degrade the accuracy from over $65\%$ to below $25\%$, with an accuracy decrease of more than $40\%$.
Breakdown on different premise orders. We present the results of fine-grained breakdown on premise ordering in Figure 5, where the orders are categorized based on Kendall tau distance $\tau$ as described in Section 2.1. Interestingly, while the top preference of all LLMs is the forward order, their preferences on other orders are not alike. Specifically, GPT-4-turbo generally prefers the backward order over other orders, and the overall performance decreases with a smaller absolute value of $\tau$. This observation is also consistent with the human reasoning pattern, as backward chaining is another well-established inference method. On the other hand, PaLM 2-L generally performs the worst with the backward order. With the decrease of $\tau$ (i.e., the premise order deviates more from the forward order), the accuracy drops. The preferences of Gemini 1.0 Pro and GPT-3.5-turbo are less consistent, still they prefer the backward order more often than other non-forward premise orders.
@@ 3.3 R-GSM for Mathematical Reasoning
Figure 7: R-GSM results with different numbers of reasoning steps in the ground truth. See Table 12 in Appendix F for accuracy numbers.
@@
Breakdown of problem complexity. Figures 7 and 8 present the breakdown results on different number of reasoning steps and different number of problem sentences, respectively. Unsurprisingly, across all LLMs, the proof accuracy suffers on problems that require more reasoning steps and contain a greater number of sentences. Overall, the gap between the accuracies on initial and rewritten problems is more significant with more reasoning steps and longer problems for both GPT-4-turbo and Gemini 1.0 Pro, while the gap remains similar across different numbers of reasoning steps and problem lengths for PaLM 2-L and GPT-3.5-turbo.
@@ 4 Related Work
Failure modes of LLMs. The premise order effect in this work is connected to several failure modes of LLMs in the literature, including the reversal curse (Berglund et al., 2023), distractibility (Shi et al., 2023), position bias (Liu et al., 2024; Wang et al., 2023), and limited capability of logical reasoning (Han et al., 2022; Xu et al., 2023; Saparov et al., 2023; Saparov and He, 2022; Wan et al., 2024; Zhu et al., 2023; Yan et al., 2023). Specifically, Shi et al. (2023) show that including irrelevant context in the problem statement leads to a considerable performance drop on GSM8K and other reasoning benchmarks, revealing that LLMs are distractible. This finding is in-line with our evaluation on logical reasoning, where we observe that adding irrelevant rules not only degrades the overall logical reasoning performance, but also escalates the premise order effect. The Reversal Curse (Berglund et al., 2023) unveils another perspective of the order effect, where they show that an LLM that recognizes “A is B” does not necessarily learn that “B is A.” While their work studies the order effect between two entities within a single factual statement, our work focuses on reasoning problems with multiple premises, without restrictions on the number of (or relationship between) entities. In particular, for logical reasoning, we demonstrate that random permutations of premises often result in worse accuracy than the purely backward order. Liu et al. (2024) discover the lost-in-the-middle phenomenon in the long-context scenario: the LLM performance is the best when the relevant information to solve the task is placed at the beginning or the end of the input context, while the performance is the worst when the LLM needs to utilize input context in the middle. In Appendix D, we show that lost-in-the-middle phenomenon does not affect the performance on our tasks, since the length of input problems does not exceed 300 tokens in our benchmark, which is relatively small compared to the context length limit of LLMs in our evaluation. Yan et al. (2023) present an approach called Concise and Organized Perception for deductive reasoning, which first generates directed graphs by connecting facts and rules in the problem, then prune and reorder the context accordingly before calling the LLM to solve the problem. The improvement achieved by this approach again demonstrates the effect of premise ordering and irrelevant premises on logical reasoning. While such input preprocessing methods can mitigate the ordering effect on certain reasoning tasks, they require task-specific design and do not generalize across domains. We consider developing generic end-to-end reasoning techniques for LLMs to address the premise order effect as future work.
@@
Order effect of language models. Some prior works show that language models are able to understand permuted texts to some extent, i.e., after a random permutation of words, models usually preserve a reasonable performance (Abdou et al., 2022; Sinha et al., 2020). Moreover,  Cao et al. (2023) show that even when a large fraction of words are scrambled, GPT-4 still achieves decent performance on several reasoning benchmarks. In contrast to permuted texts in these works that are typically unnatural and nonsensical, our premise order permutations do not alter the semantic meaning and remain syntactically valid (we manually verify this). Nevertheless, we demonstrate that LLM reasoning performance is highly brittle to the ordering of the premises. For long-digit addition, prior works demonstrate that reversing the input numbers is a key to achieve better length generalization performance (Lee et al., 2023; Zhou et al., 2023, 2024). Specifically, by reversing the input numbers so that the least significant digit is presented first, the Transformer learns a simpler way of performing addition, where the model only needs to perform computation with the corresponding digits of operands and the carry-on digit at each step, without the need of looking at other digits. This approach enables the Transformer to better perform addition when trained from scratch, which also aligns with our finding: after reversing the input numbers, the premise order (i.e., orders of digits) follows the right ordering of performing long-digit addition, thus enables Transformers to better learn the task.
@@ Appendix C R-GSM Examples
Appendix D Does Logical Reasoning Suffer from the Lost-in-the-middle Issue?
Liu et al. (2024) demonstrate that when the input context becomes long, LLMs might suffer from the lost-in-the-middle issue: the model performance significantly degrades when relevant information to solve the task is in the middle of the input, instead of at the beginning or the end. Therefore, when given distracting rules for logical reasoning, another potential factor that might affect the model performance is the position of relevant rules in the model input.
To examine the effect of such position bias, we conduct ablations on PaLM 2-L with 10 distracting rules, and we compare the performance with relevant rules added in the beginning, middle or the end of the problem description. Table 5 shows that with the same order and number of rules, the variation in performance is very small, whereas changing the order significantly affects the results. Note that the longest inputs in our logical reasoning benchmark, i.e., problems with 12 relevant rules and 10 distracting rules, only contain no more than 300 tokens, which is relatively short compared to the context length limit of LLMs in our evaluation. These results confirm that on our tasks where the input problems (and thus input context) are short, lost-in-the-middle phenomenon is not the primary cause of the performance difference. In our primary experiments, for all logical reasoning problems, we interleave distracting rules with relevant rules in the input context.
[…]
8, beginning 68.0% 40.0% 45.5%
8, middle 67.0% 39.0% 44.5%
8, end 67.0% 40.0% 45.5%
12, beginning 36.5% 17.0% 16.0%
12, middle 36.5% 17.0% 18.5%
12, end 35.0% 16.0% 19.5%
Table 5: Logical reasoning results performance of PaLM 2-L, with relevant rules at different positions of the input context.
Appendix E Full Results for Logical Reasoning
Tables 6 and 9 present the accuracy numbers for Figures 3 and 5, which are results on different numbers of relevant rules without distracting rules.
Tables 7 and 10 present the accuracy numbers for Figures 4 and 6 with 5 distracting rules.
Tables 8 and 11 present the accuracy numbers for Figures 4 and 6 with 10 distracting rules.
@@ 6 Forward 100%
(a) GPT-4-turbo.
# Rules Order Acc
4 Forward 98.5%
Backward 98.5%
Shuffled 98.3%
5 Forward 98.5%
Backward 98.5%
Shuffled 98.3%
6 Forward 98.0%
Backward 93.5%
Shuffled 95.3%
7 Forward 96.5%
Backward 89.0%
Shuffled 91.2%
8 Forward 95.5%
Backward 77.0%
Shuffled 87.7%
9 Forward 94.0%
Backward 79.0%
Shuffled 85.7%
10 Forward 95.0%
Backward 75.5%
Shuffled 81.0%
11 Forward 94.0%
Backward 66.0%
Shuffled 78.7%
12 Forward 88.0%
Backward 57.5%
Shuffled 66.5%
@@
Appendix F Full Results on R-GSM
Tables 12 and 13 present the accuracy numbers for Figures 7 and 8, which are breakdown results on R-GSM problems with different numbers of reasoning steps and different numbers of sentences in the problem description respectively.